OpenStudy (anonymous):

Write an equation in standard form for the ellipse with center (0, 0), vertex at (6, 0) and co-vertex at (0, 4).

OpenStudy (anonymous):

I am a first-time college math student and I don't understand any of it and it to be broken down for me. Thanks

OpenStudy (anonymous):

your welcome?

OpenStudy (amistre64):

\[\frac{x^2}{36}+\frac{y^2}{16}=1\]

OpenStudy (anonymous):

thx amistre64 always coming to my rescue :)